OpenStudy (anonymous):

how do i set up the equation for this? Jordan is a manager of a car dealership. He has 3 professional car washers to clean the entire lot of cars, Jennifer, Arianna, and Matthew. Jennifer can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Matthew is new to the car dealership, so no one knows how long it will take him. Jordan assigns all of them to wash the cars together. Explain to Jordan how this task can tell him how long it would take Matthew to complete the task if he worked by himself. Use complete sentences.

OpenStudy (anonymous):

@Hannah_Waller

OpenStudy (anonymous):

give me a medal for helping you last time and ill tell you

OpenStudy (anonymous):

OK where do you think you start?

OpenStudy (anonymous):

@Hannah_Waller I know the equation up to this point 1/14 + 1/11 + 1/M = 1/t

OpenStudy (anonymous):

@bigsexy75 nice way to help someone out.

OpenStudy (anonymous):

then you end up getting \[25 \over 154\]

OpenStudy (anonymous):

So you know this, Has a total of 3 car washers J- Can wash cars in 14 hours A- Can wash cars in 11 hours M- ?

OpenStudy (anonymous):

so im not sure how you got decimas... im confuseed haha

OpenStudy (anonymous):

decimals?

OpenStudy (anonymous):

haha i meant fractions

OpenStudy (anonymous):

Their not fractions, more like proportions

OpenStudy (anonymous):

that's just how you set up the problem

OpenStudy (anonymous):

oh ok sorry i dont completly get it... :(

OpenStudy (anonymous):

@phi could you please help understand this?

OpenStudy (phi):

I guess you have to solve for M

OpenStudy (anonymous):

Yes, but I don't know how to get it that

OpenStudy (phi):

one technique is to "clear the denominators" by multiply the equation by 11*14*t (all terms both sides)

OpenStudy (anonymous):

wait, what do you mean? sorry, this is really confusing for me

OpenStudy (phi):

like this \[ 14\cdot 11\cdot t \left( \frac{1}{11} +\frac{1}{14} +\frac{1}{m} \right) = 14\cdot 11\cdot t \cdot \frac{1}{t} \]

OpenStudy (anonymous):

ohh okay, so once i've got that, what do i do?

OpenStudy (phi):

I have not done this, so you can work along and see how it goes.

OpenStudy (anonymous):

I've never done it this way. I usually just cross multiply and then divide to get my answer, but this one includes matthew so it's confusing

OpenStudy (phi):

yes, 3 things is more complicated. but distribute the 11*14*t what do you get ?

OpenStudy (phi):

if you don't like this way, we can do it using fractions starting with \[ \frac{1}{11} +\frac{1}{14} +\frac{1}{m} = \frac{1}{t} \] add the first two numbers (which you did up above) to get 25/154 \[ \frac{25}{154} +\frac{1}{m} = \frac{1}{t} \] subtract 25/154 from both sides \[ \frac{25}{154} - \frac{25}{154} +\frac{1}{m} = \frac{1}{t} -\frac{25}{154}\\ \frac{1}{m} = \frac{1}{t} -\frac{25}{154} \]

OpenStudy (phi):

now combine the left side using a common denominator of 154 t \[ \frac{1}{m} = \frac{1}{t} -\frac{25}{154} \\ \frac{1}{m} = \frac{1}{t}\cdot \frac{154}{154} -\frac{25}{154}\cdot \frac{t}{t} \\ \frac{1}{m} = \frac{154-25t}{154t} \]

OpenStudy (phi):

are you following?

OpenStudy (phi):

*now combine the *right* side using a common denominator of 154 t

OpenStudy (anonymous):

that's where i'm confused, what exactly am i combining?

OpenStudy (phi):

\[ \frac{1}{t}- \frac{25}{154} \]

OpenStudy (phi):

do you see you want a common denominator of 154t ?

OpenStudy (anonymous):

ohhh okay, sorry i wasn't looking at the right thing, yes i see that

OpenStudy (phi):

so multiply the first fraction by 154/154 and the second fraction by t/t

OpenStudy (anonymous):

yes, got that

OpenStudy (phi):

did you get all the steps so far?

OpenStudy (anonymous):

Yes I do, now for the last part do we have to multiply 154t on both sides?

OpenStudy (phi):

we now have \[ \frac{1}{m} = \frac{154-25t}{154t} \] the simplest thing to do is "flip" (i.e. invert) both sides \[ m = \frac{154t}{154-25t} \]

OpenStudy (anonymous):

okay, i don't know what to do from there

OpenStudy (phi):

Explain to Jordan how this task can tell him how long it would take Matthew to complete the task alone

OpenStudy (phi):

you tell Jordan to time how long all 3 take to do the job. That will be "t" in your equation put in that number for "t", and solve (figure out) how fast M can do the job by himself.

OpenStudy (phi):

For example, if it takes all 3 people 3 hours to do the job, then M= 3*154/(154-3*25) = 5.85 hours or about 5 hours 51 minutes

OpenStudy (anonymous):

so whatever number i get i plug it in for t and that is my answer

OpenStudy (phi):

Yes. The question is asking how to figure out how long M takes working alone. You explain how by saying, time how long all 3 take to do the job. Use that time in the equation to find M

OpenStudy (phi):

You can't find M unless you know how long it takes the 3 to do the job, and we don't know that.

OpenStudy (anonymous):

oh okay, got it. Thank you so much for your help and patience, i appreciate it!! :)

OpenStudy (phi):

so you get it that we don't know "t"? right? but if we did, we know how to find M.

OpenStudy (anonymous):

yes, i got that